# Participation bias in the estimation of heritability and genetic correlation

### v2

@@ 3 Discussion
In the last decade, many methods have been proposed to estimate heritability and genetic correlation. Despite the success, estimates are typically derived under the assumption of random sampling, without taking PB into account. In addition, the non-random component underlying participation is often implicitly thought of as a function of environmental factors and established phenotypes such as educational attainment, frequency of alcohol use, etc. Following this line of reasoning, to adjust for PB, a recent study15 applied IPW to the UKBB data, with a propensity score constructed from other phenotypes. Apart from the limitations that the propensity score was limited by the availability of phenotypes that could be harmonized between UKBB and other random sampling studies, most importantly, the propensity score does not include genotypes. For this IPW adjustment to be sufficient for analyses that include genotypes, the main focus of genetic studies that include heritability and genetic correlation estimates, it requires that the genetic component underlying participation manifests its effect entirely through the propensity score. Under this assumption, the participation polygenic score constructed in ref.16 (denoted as pPGS), based on a GWAS that did not use any information on other phenotypes, should not have any predictive power on other phenotypes conditioning on the propensity score. We examined this with four phenotypes, EA, BMI, and the invitation and participation in the secondary physical activity study. We observed significant associations ($p<0.05/4$, Bonferroni correction) between pPGS and all four phenotypes before and after adjustment for the predicted participation indicated by the propensity score (Table 4), indicating that the assumption does not hold. Similar analyses were performed in ref.16 using EA in the place of the propensity score. These results suggest that participation should be treated as a complex behavioral trait in its own right with its specific genetic component, and not simply a consequence of other established phenotypes.
Table 4: The pPGSs association with phenotypes adjusted for predicted participation. Linear regression was used for analyzing EA and BMI, and logistic regression was used for the two secondary invitation and participation for physical activity study. Displayed are fitted coefficients (Effect/log(OR)) and $p$-values from regressing phenotypes on the pPGS in a subset of White British unrelateds, taking sex, year of birth, age at recruitment, and 40 principal components as covariates. The pPGS and the quantitative phenotypes were transformed to have a variance of 1 and thus effect is in s.d. units.
Phenotypes pPGS pPGS adjusted for predicted participation
Effect $p$-value Sample size Effect $p$-value Sample size
EA $3.1\times 10^{-2}$ $<2\times 10^{-16}$ 268,205 $2.0\times 10^{-2}$ $<2\times 10^{-16}$ 227,688
BMI $-1.9\times 10^{-2}$ $<2\times 10^{-16}$ 270,330 $-1.6\times 10^{-2}$ $8.9\times 10^{-15}$ 227,688
Secondary invitation $2.7\times 10^{-2}$ $3.5\times 10^{-12}$ 132,148/139,048 $2.1\times 10^{-2}$ $9.0\times 10^{-7}$ 114,834/112,854
Secondary participation $3.0\times 10^{-2}$ $7.6\times 10^{-8}$ 59,264/72,884 $2.6\times 10^{-2}$ $1.7\times 10^{-5}$ 52,317/62,517
In this article, we build a statistical model for understanding the influence of PB on estimating heritability and genetic correlation. We regard participation as a phenotype in its own right, which can be affected by a combination of genetic and non-genetic factors. Compared with the weighting strategy, our model utilizes the IBD-based information and provides extra degrees of freedom. We derived theoretical results for the expected differences between the heritability and genetic correlation estimates obtained from the biased sample and the actual values of the population. We directly model the GWAS summary statistics in the selected sample, without using the genotypic information of the non-participants. We validated the theoretical results with simulations and applied our results to 12 UKBB phenotypes. We found 8 of the phenotypes were significantly associated with participation of the UKBB. We also found that, without adjustments, the heritability and the absolute value of the genetic correlation estimates had underestimation bias in the sample of participants.
@@
There are several limitations to the current approach. First, the IBD-based participation GWAS we used only captured direct effects, and we did not consider the indirect effects of participation. Second, the sample size of IBD siblings for deriving participation statistics is relatively limited. Although we have demonstrated via both simulations and real data applications that the standard errors before and after our adjustment are comparable, larger sample sizes will lead to more accurate results with smaller MSEs. Third, in order to adjust for PB, we leveraged the HSE dataset to derive the mean shift between UKBB and HSE. We assumed that the HSE dataset was based on random sampling, as it has incorporated weighting to account for nonresponse bias. Violation of this assumption would reduce the effectiveness of the adjustments, but most likely they would be in the right direction. Lastly, the adjustments were calculated assuming the invited list, about 9.5 million in size for the UKBB, is representative of the target population, about 21 million in size with the age constraint 13. Violation of this assumption does not affect the estimate of the genetic component underlying the overall participation bias that incorporates both invitation bias and agreeing to participate if invited bias. However, the estimate of the strength of this genetic component, $e.g.$ its heritability, could be impacted, and through that affect the other adjustments. The exact effect is mathematically complicated and depends on many factors, but we can get an idea by considering two extreme scenarios for the UKBB study. One extreme is that all those who were not invited would not have agreed to participate if invited anyway. In this case, the appropriate adjustments would be smaller than what are currently estimated, $e.g.$ the adjustment of heritability estimates for other phenotypes should be shrunk by about $25\%$. The other extreme is if all the bias came from the invitation list, and participation upon invitation was completely random. In that case, our current adjustments would be in the right direction but conservative, $i.e.$ smaller than what should be. Given that, we believe the adjustments we provided are reasonable. To do substantially better, more complicated modelling and additional information, such as phenotypic differences between target population and the invited list, are probably required.
@@
Despite the limitations, our model and method provide insights about the genetic architecture underlying participation and other phenotypes. It is, we believe, an important first step towards adjusting PB in genetic studies in a way that does not rely entirely on phenotypic differences between sample and population. For future research, it is conceptually advantageous to develop statistical methods to investigate the underlying causality of participation and other phenotypes. In particular, it is often assumed that EA has a causal effect on participation 15, while whether and to what extent the inclination to participate, as a behavioral phenotype, has a causal impact on other phenotypes including EA remains unclear. A deeper understanding of causality will be helpful in understanding the underlying mechanisms and dynamics of human behavior.
@@ 4 Methods
4.1 Simulation settings
We simulated a population with $5\times 10^{6}$ sibling pairs and $10^{4}$ SNPs. The genotypes were generated from a binomial distribution with block-wise autoregressive (AR(1)) LD structure using the R package CorBin 19. Each block has $50$ SNPs, with the correlation matrix $\bm{R}_{l}$ of the $l$-th LD block as
$\mathbf{R}_{l}={\left(\begin{array}[]{cccc}1&\rho_{l}&\cdots&\rho_{l}^{p_{l}-1}\\ \rho_{l}&1&\cdots&\rho_{l}^{p_{l}-2}\\ \vdots&\vdots&\ddots&\vdots\\ \rho_{l}^{p_{l}-1}&\rho_{l}^{p_{l}-2}&\cdots&1\end{array}\right)},$ (1)
where $p_{l}$ is the number of SNPs in the block, and $\rho_{l}\sim$Unif$(0.1,0.9)$. The allele frequencies in each block were also sampled from Unif$(0.1,0.9)$. The correlations are higher for adjacent variants and decrease with the increasing distance between the variants, which mimics the real LD structures 20. We set $10\%$ of the SNPs to have nonzero effects. Specifically, for the $j$-th SNP with nonzero effects, $\beta_{j}$ and $\gamma_{j}$ were simulated from a bivariate normal distribution with mean 00 and variance $h^{2}_{x}/1000$ and $h^{2}_{y}/1000$, with correlation $\rho_{g}$. The non-genetic terms were also simulated from a bivariate normal distribution with mean 00, variance ($1-h^{2}_{x}$) and ($1-h^{2}_{y}$), and correlation $\rho_{e}$. For each time of the simulation, we generated the liability score of participation for each sample with $X=\boldsymbol{G}\boldsymbol{\beta}+{\epsilon}_{x}$, and another phenotype with $Y=\boldsymbol{G}\boldsymbol{\gamma}+{\epsilon}_{y}$, where $\boldsymbol{G}$ is the standardized genotype. As for the analysis of two phenotypes, we generated the effect sizes of both phenotypes, as well as the effect sizes of participation liability score with a 3-dimensional normal distribution. We set the participation rate ($\alpha$) and sibling recurrence risk ($\lambda$) to $0.055$ and $2$, which equals to the corresponding values reported in the UKBB 7; 21; 16. Samples with $X\geq\Phi^{-1}(1-\alpha)$ were selected as participants. We only included one of the sibling pairs, so the sample size of participants was around $0005\times 10^{6}\times 0.055=275,$ 000. We then performed GWAS on $Y$ in the sample of participants. We followed the procedure in Benonisdottir and Kong (2023) to derive the test statistics of participation with IBD-based between-sib-pairs comparison information. Simulations were repeated 50 times.
@@ 4.2 UKBB summary statistics
The GWAS summary statistics of UKBB participation were directly downloaded from GWAS catalog 16. Other UKBB summary statistics based on the genetic data of 274,485 white British unrelated individuals in the UKBB after quality control 21. The phenotypes of interest were adjusted for year of birth (data-field 34), age at recruitment (data-field 21022), and sex (data-field 31) when applicable. We further adjusted top 40 principal components 21. Quantitative phenotypes except for EA, ALC, and WLP, were rank-based inverse normal transformed 22 separately for each sex. A detailed description of the process of phenotypes is provided in Supplementary Note 1.6. We used the same quality filtering protocol for the sequence variants as Benonisdottir and Kong (2023) 16. The analysis was restricted to the set of 500,632 high-quality variants with the missing rate $<5\%$ and with MAF $>1\%$. We used PLINK 1.90 to derive GWAS summary statistics 23.
@@ 4.3 HSE datasets
We obtained the anthropometric measures from 81,118 individuals from the Health Survey for England (HSE) for the years 2006-2010 24; 25; 26; 27; 28, which consists of an annual cross-sectional survey. The samples are a representative population of England through a two-stage random probability sampling process 13. The HSE data have incorporated weighting to account for nonresponse bias since 2003 29. A detailed description of the collection of HSE datasets is in Supplementary Note 1.6. For the computation of the mean shift between HSE and UKBB, we restricted samples of white British ancestry and ages from 40 to 65, resulting in 20,208 individuals. The phenotypes in both HSE and UKBB were adjusted for sex, age, age2, sex*age, and sex*age2.
@@ 4.4 LD score regression
We used LD score regression (LDSC, v.1.0.1) to derive the estimates for heritability and genetic correlation, which ignores the effects of PB. In the simulations, the LD scores were computed with true LD matrices. The LD scores in real data analyses were computed by the Pan-UKB team 30 (downloaded on 7 April 2021).
@@ 4.5 Polygenic score analysis
We computed the pPGS with PLINK 1.90 23, which summed over the weighted genotypes of the $500,632$ SNPs after quality control. The $z$ scores were used as weights, which were transformed from the combined $p$-values in the GWAS summary statistics of participation provided in ref. 16. The pPGS was standardized to have variance 1, and the relationship between the pPGS and EA, BMI, secondary invitation (binary), and secondary participation (binary) was estimated with a linear regression and logit regression in R (v.4.3.1), in the group of White British unrelateds. The sex, year of birth, age at recruitment, genotyping array, and 40 principal components were used as covariates.
@@ 6 Code availability
The replication codes are available at https://github.com/shuangsong0110/ParticipationBias.

### v3

@@ 3 Discussion
In the last decade, many methods have been proposed to estimate heritability and genetic correlation. Despite the success, estimates are typically derived under the assumption of random sampling, without taking PB into account. In addition, the non-random component underlying participation is often implicitly thought of as a function of environmental factors and established phenotypes such as EA, frequency of alcohol use, etc. Following this line of reasoning, to adjust for PB, a recent study15 applied IPW to the UKBB data, with a propensity score constructed from other phenotypes. Apart from the limitations that the propensity score was limited by the availability of phenotypes that could be harmonized between UKBB and other random sampling studies, most importantly, the propensity score does not include genotypes. For this IPW adjustment to be sufficient for analyses that include genotypes, the main focus of genetic studies that include heritability and genetic correlation estimates, it requires that the genetic component underlying participation manifests its effect entirely through the propensity score. Under this assumption, the participation polygenic score constructed in Benonisdottir and Kong (2023) 16 (denoted as pPGS), based on a GWAS that did not use any information on other phenotypes, should not have any predictive power on other phenotypes conditioning on the propensity score. We examined this with four variables that are significantly associated with the pPGS: EA, BMI, and the invitation and participation in a physical activity study. When adjusting for the propensity score constructed by Schoeler et al (2023), the associations between the pPGS and the four variables shrunk by some degree, but remained highly statistically significant ($p$ ranges from $1.7\times 10^{-5}$ to $<2\times 10^{-16}$, see Supplementary Note 1.8 and Supplementary Table 2), indicating that the assumption does not hold. Similar analyses were performed in Benonisdottir and Kong (2023) 16 using EA in the place of the propensity score. These results support the belief that participation should be treated as a complex behavioral trait in its own right with its specific genetic component, and not simply a consequence of other established phenotypes.
In this article, we build a statistical model for understanding the influence on estimates of parameters involving phenotypes of interest. Conceptually, to properly adjust estimates of parameters that involve both genotypes and phenotypes, $e.g.$ heritability and genetic correlation, information on the differences between sample and population for both genotypes and phenotypes is necessary. For genotype differences, we utilize the IBD-based information on the genetic component of participation 16, which distinguishes our approach from existing methods such as IPW. For phenotype differences, we utilize the mean shifts from population to sample of the phenotypes of interest. Further research would be to explore what other information, if available, could be used to improve the adjustments and allows us to entertain more complex PB models. We applied our adjustment method to 12 UKBB phenotypes. We found 8 of the phenotypes were significantly associated with participation of the UKBB. We also found that, without adjustments, the heritability and the absolute value of the genetic correlation estimates had underestimation bias in the sample of participants. We note that LDSC estimates SNP-based heritability, $i.e.$ the proportion of phenotypic variance attributable to a given set of SNPs20; 21. This would increase as the number of SNPs increases and is in general a fraction of the heritability. Assuming the genetic architecture of the SNP set is representative of the entire genome, we believe the expected value of the ratio between the adjusted and unadjusted estimates should be stable as the number of SNPs increases. As to the adjusted genetic correlation estimates, they should be estimates for the full genetic components also.
@@
There are several limitations to the current approach. First, the IBD-based participation GWAS we used only captured direct effects, and we did not consider the indirect effects of participation. Second, the sample size of IBD siblings for deriving participation statistics is relatively limited. Although we have demonstrated via both simulations and real data applications that the standard errors before and after our adjustment are comparable, larger sample sizes will lead to more accurate results with smaller MSEs. Third, in order to adjust for PB, we leveraged the HSE dataset to derive the mean shift between UKBB and HSE. We assumed that the HSE dataset was based on random sampling, as it has incorporated weighting to account for nonresponse bias. Violation of this assumption would reduce the effectiveness of the adjustments, but most likely they would still be in the right direction. Lastly, the adjustments were calculated assuming the invited list, about 9.5 million in size for the UKBB, is representative of the target population, about 21 million in size with the age constraint 13. Violation of this assumption does not affect the estimate of the genetic component underlying the overall PB that incorporates both invitation bias and agreeing to participate if invited bias. However, the estimate of the strength of this genetic component, $e.g.$ its heritability, could be impacted, and through that affect the other adjustments. The exact effect is mathematically complicated and depends on many factors, but we can get an idea by considering two extreme scenarios for the UKBB study. One extreme is that all those who were not invited would not have agreed to participate if invited anyway. In this case, the appropriate adjustments would be smaller than what are currently estimated, $e.g.$ the adjustment of heritability estimates for other phenotypes should be shrunk by about $25\%$. The other extreme is if all the bias came from the invitation list, and participation upon invitation was completely random. In that case, our current adjustments would be in the right direction but conservative, $i.e.$ smaller than what should be. Given that, we believe the adjustments we provided are reasonable. To do substantially better, more complicated modelling and additional information, such as phenotypic differences between target population and the invited list, are probably required.
@@
Despite the limitations, our model and method provide insights about the genetic architecture underlying participation and other phenotypes. It is, we believe, an important first step towards adjusting PB in genetic studies in a way that does not rely entirely on phenotypic differences between sample and population. For future research, it is conceptually advantageous to develop statistical methods to investigate the underlying causality of participation and other phenotypes. In particular, it is often assumed that EA has a causal effect on participation 15. By contrast, whether and to what extent the inclination to participate, as a behavioral phenotype, has a causal impact on other phenotypes including EA has not been seriously considered. A deeper understanding of causality will be helpful in understanding the underlying mechanisms and dynamics of human behavior.
@@ 4.1 UKBB summary statistics
The GWAS summary statistics of UKBB participation were directly downloaded from GWAS catalog 16. Other UKBB summary statistics based on the genetic data of 274,485 white British unrelated individuals in the UKBB after quality control 22. The phenotypes of interest were adjusted for year of birth (data-field 34), age at recruitment (data-field 21022), and sex (data-field 31) when applicable. We further adjusted top 40 principal components 22. Quantitative phenotypes except for EA, ALC, and WLP, were rank-based inverse normal transformed 23 separately for each sex. A detailed description of the process of phenotypes is provided in Supplementary Note 1.6. We used the same quality filtering protocol for the sequence variants as Benonisdottir and Kong (2023) 16. The analysis was restricted to the set of 500,632 high-quality variants with the missing rates $<5\%$ and with minor allele frequencies $>1\%$. We used PLINK 1.90 to derive GWAS summary statistics 24.
@@ 4.2 HSE datasets
We obtained the anthropometric measures from 81,118 individuals from the Health Survey for England (HSE) for the years 2006-2010 25; 26; 27; 28; 29, which consists of an annual cross-sectional survey. The samples are a representative population of England through a two-stage random probability sampling process 13. The HSE data have incorporated weighting to account for nonresponse bias since 2003 30. A detailed description of the collection of HSE datasets is in Supplementary Note 1.6. For the computation of the mean shift between HSE and UKBB, we restricted samples of white British ancestry and ages from 40 to 65, resulting in 20,208 individuals. The phenotypes in both HSE and UKBB were adjusted for sex, age, age2, sex*age, and sex*age2.
@@ 4.3 LD score regression
We used LD score regression (LDSC, v.1.0.1) to derive the estimates for heritability and genetic correlation, which ignores the effects of PB. In the simulations, the LD scores were computed with true LD matrices. The LD scores in real data analyses were computed by the Pan-UKB team 31 (downloaded on 7 April 2021).
@@ 4.4 Polygenic score analysis
We computed the pPGS with PLINK 1.90 24, which summed over the weighted genotypes of the $500,632$ SNPs after quality control. The $z$ scores were used as weights, which were transformed from the combined $p$-values in the GWAS summary statistics of participation provided in Benonisdottir and Kong (2023) 16. The pPGS was standardized to have variance 1, and the relationship between the pPGS and EA, BMI, secondary invitation (binary), and secondary participation (binary) was estimated with a linear regression and logit regression in R (v.4.3.1), in the group of White British unrelateds. The sex, year of birth, age at recruitment, genotyping array, and 40 principal components were used as covariates.
@@ 6 Code availability
We have developed an R package for adjusting the effects of PB on the estimation of heritability and genetic correlations, which is available at https://github.com/shuangsong0110/ParticipationBias.